0 0
Add a comment Improve this question Transcribed image text
Answer #1
Concepts and reason

The concept used to solve this problem is based on molecular orbital diagram.

A molecular orbital diagram is used to explain chemical bonding in a molecule. This diagram is based on the molecular orbital theory.

Fundamentals

According to molecular orbital theory molecular orbitals of a molecule is formed by combination of its atomic orbitals and electrons present in molecule are distributed among the molecular orbitals.

The total number of electrons present in O2{{\rm{O}}_2} molecule are 16. Thus, the electronic configuration for combination of two atomic orbitals of oxygen is as follows.

(σ1s)2(σ1s)2(σ2s)2(σ2s)2(σ2pz)2(π2px)2(π2py)2(π2px)1(π2py)1{\left( {{\sigma _{1s}}} \right)^2}{\left( {\sigma _{1s}^ * } \right)^2}{\left( {{\sigma _{2s}}} \right)^2}{\left( {\sigma _{2s}^ * } \right)^2}{\left( {{\sigma _{2{p_z}}}} \right)^2}{\left( {{\pi _{2{p_x}}}} \right)^2}{\left( {{\pi _{2{p_y}}}} \right)^2}{\left( {\pi _{2{p_x}}^ * } \right)^1}{\left( {\pi _{2{p_y}}^ * } \right)^1}

Here sigma (σ)\left( \sigma \right) and pi (π)\left( \pi \right) orbitals are called bonding molecular orbitals. The sigma star (σ)\left( {{\sigma ^*}} \right) and pi star (π)\left( {{\pi ^*}} \right) are called antibonding orbitals.

In O2{{\rm{O}}_2} orbital, the valence orbitals are 2s and 2p and total number of electrons in these orbitals are 12. Thus, Construct valence molecular orbital diagram and fill 12 electrons as follow.

Ans: